OpenStudy (anonymous):

Mandy loves to eat New York pizza. Each square inch of pizza is 10 calories. She orders a slice of pizza from a pizza with a diameter of 20 inches that is cut with a central angle of 50 degrees. How many calories does she eat? about 314 calories about 436 calories about 3,142 calories about 43 calories

OpenStudy (pfenn1):

The area of the pizza slice can be determined by the area of the whole circle multiplied by the percentage of the circle, the slice takes up.

OpenStudy (anonymous):

Area of a circle is \[\pi r^2\] Radius is 10 inches. If it is cut with 50 degree angle, then 50/360 is eaten. Find the area of the circle: \[\pi (10)^2\] Multiply by \[\frac{50}{360}\] You will get around 43.

OpenStudy (anonymous):

its not 43

OpenStudy (anonymous):

Oops sorry. Forgot to account for the fact that each square inch is 10 calories. So multiply that by 10, or 436.

Directrix (directrix):

I got 436.33 but need to check my work. Especially after I missed the first problem in this series.